# Human-in-the-Loop Feature Selection Using Interpretable Kolmogorov-Arnold Network-based Double Deep Q-Network

### v1

@@ V-D2 Visualization
To understand the importance of features, we utilize the plotting functionality of the KAN model. Fig. 3a visualizes the activation functions of the neurons, and Fig. 3c shows the training steps of the KAN agent where the bottom layer indicates 64 input variables, followed by 8 hidden neurons and then 10 outputs. The transparency of each activation function $\phi_{l,i,j}$ is set proportionally to $\tanh(\beta A_{l,i,j})$, where $\beta=30$. Smaller $\beta$ allows us to focus on more significant activations.
V-D3 Symbolification
[…]
VI Conclusion and Future Work
@@
This study introduced a HITL feature selection framework using a DDQN architecture with KAN, achieving interpretable, per-example feature selection and improved performance across multiple metrics. Through simulated feedback and stochastic feature sampling from diverse distributions, the HITL-KAN-DDQN model dynamically selects feature subsets for each observation, focusing on the most relevant features and improving interpretability. Operating on low-resolution $8\times 8$ pixel images, our model consistently outperforms MLP-based DDQN models, achieving higher accuracy while using fewer neurons—specifically, an 8-neuron single hidden layer compared to the 32-neuron layer required by the MLP-DDQN. This compact design provided computational efficiency without compromising predictive power, confirming the effectiveness of the HITL-KAN-DDQN model for feature selection.
Future work could improve simulated feedback to align better with expert annotations and apply this HITL approach to various feature selection tasks and high-dimensional domains. Exploring RL further may involve querying feedback based on model uncertainty or data point importance, moving toward active learning. Other RL policies to explore include DQN, REINFORCE, proximal policy optimization (PPO), advantage actor-critic (A2C), and soft actor-critic (SAC).

### v2

@@ 5.7.2 Visualization
To understand the importance of features, we utilize the plotting functionality of the KAN model. Fig. 4a visualizes the activation functions of the neurons, and Fig. 4c shows the training steps of the KAN agent, where the bottom layer indicates 64 input variables, followed by 8 hidden neurons, and then 10 outputs. The transparency of each activation function $\phi_{l,i,j}$ is set proportionally to $\tanh(\beta A_{l,i,j})$, where $\beta=30$. Smaller $\beta$ allows us to focus on more significant activations.
5.7.3 Symbolification
Our approach identifies symbolic forms within the KAN architecture to enhance interpretability. We sample network preactivations $x$ and post-activations $y$, fitting an affine transformation $y\approx cf(ax+b)+d$, where $a$, $b$ adjust inputs, and $c$, $d$ scale and shift outputs. This fitting involves grid search on $a$, $b$ and linear regression for $c$, $d$. Using a library of symbolic functions ($x,x^{2},x^{3},\exp(x),\log(x),\sqrt{x},\sin(x),|x|$), auto-symbolic regression replaces learned activations with interpretable formulas. The policy in KAN-DDQN is represented as $a_{i}=f_{i}(x)$ for each output $i$, with the overall action $a=\underset{i}{\mathrm{argmax}}\ a_{i}$. Fig. 4b displays the importance of features extracted from a symbolic formula, showing their frequency of occurrence and corresponding contributions to the model’s predictions. Fig. 5 illustrates the activation function of a middle neuron in the KAN’s 8-neuron hidden layer for two input neurons, with the x-axis showing input values and the y-axis indicating the activation output after the spline function.
6 Conclusion and Future Work
@@
This study introduced a HITL feature selection framework using a DDQN architecture with KAN, achieving interpretable, per-example feature selection and improved performance across multiple metrics. Through simulated feedback and stochastic feature sampling from diverse distributions, the HITL-KAN-DDQN model dynamically selects feature subsets for each observation, focusing on the most relevant features and improving interpretability. Operating on low-resolution $8\times 8$ pixel images, our model consistently outperforms MLP-based DDQN models, achieving higher accuracy while using fewer neurons, specifically, an 8-neuron single hidden layer compared to the 32-neuron layer required by the MLP-DDQN. This compact design provided computational efficiency without compromising predictive power, confirming the effectiveness of the HITL-KAN-DDQN model for feature selection.
Future work will extend this HITL feature selection framework beyond low-resolution vision benchmarks to include standard-resolution images and non-vision modalities such as tabular and multimodal datasets. This will enable rigorous evaluation of generalization across diverse, real-world scenarios and characterize the benefits and limitations of instance-wise selection under higher-dimensional and noisier conditions. We also aim to improve simulated feedback mechanisms to better align with expert annotations, further strengthening interpretability and fairness. To address the computational demands of KAN models, future implementations will incorporate more efficient variants such as FastKAN [10] and PowerMLP [12], leveraging their lightweight architectures to reduce training time while maintaining interpretability and performance. Another critical direction is to replace simulated feedback with actual human feedback from domain experts in future iterations. By involving real experts during the interactive feature selection process, we can increase the practical relevance, interpretability, and fairness of the selected feature subsets. Integrating expert-driven annotations will also facilitate comparative analysis between simulated and authentic feedback systems, enabling refined alignment strategies. From a reinforcement learning perspective, future research will explore active learning-based query strategies that prioritize feedback acquisition based on model uncertainty and data point importance. Moreover, we intend to experiment with a broader range of reinforcement learning algorithms beyond DDQN, including DQN, REINFORCE, proximal policy optimization (PPO), advantage actor-critic (A2C), and soft actor-critic (SAC), to evaluate their suitability for adaptive, per-example feature selection in human-in-the-loop settings.